A sports car accelerates from rest to 26.8 m/s (roughlv 60 min) in 5.1 seconds. What le the acceleration the car


1
Expert's answer
2022-02-15T09:58:19-0500

Given:

"v_i=0"

"v_f=\rm 26.8\: m/s"

"\Delta t=\rm 5.1\: s"


The acceleration the car

"a=\frac{\Delta v}{\Delta t}\\
a=\rm\frac{26.8\: m/s-0\: m/s}{5.1\: s}=5.25\; m/s^2"